Check my work The manager at a water park constructed the following frequency distribution to summarize attendance in July and August Frequency 15 15 18 1,000 up to 1,250 1,250 up to 1,500 1,500 up to 1,750 1,750 up to 2,000 2,000 up to 2,250 2,250 up to 2,500 19 20 a. Construct the corresponding relative frequency, cumulative frequency, and cumulative relative frequency distributions. (Round Relative Frequency and Cumulative Relative Frequency answers to 3 decimal places.) Relative Cumulative Relative Frequency Cumulative Attendance 1,000 up to 1,250 1,250 up to 1,500 1,500 up to 1,750 1,750 up to 2,000 2,000 up to 2,250 2,250 up to 2,500 b-1. What is the most likely attendance range? (Rou nd percentage answer to the nearest whole number.) with a % frequency. The most likely attendance range is from up to Prev 28 of 30 Next>
media%2Fd4d%2Fd4d26001-8b27-4305-b086-cb
media%2Fbf0%2Fbf091be1-8e76-4ca1-988d-36
Answer #1

a) Tha Calculation table is given below:

Attendance Freq. Rel. Freq. Cum. Freq. Cum. Rel Freq.
1000 up to 1250 15 0.160 15 0.160
1250 up to 1500 15 0.160 30 0.319
1500 up to 1750 18 0.191 48 0.511
1750 up to 2000 7 0.074 55 0.585
2000 up to 2250 19 0.202 74 0.787
2250 up to 2500 20 0.213 94 1.000

b-1) The most likely attendance range is from 2250 to 2500 with a 21% frequency

b-2) Number of times= 74

c-1) Percentage of time = 16%

c-2) Percentage of time = 51%

c-3) Percentage of time = 49%